OpenStudy (anonymous):

The path of a soccer ball is modeled by the function h(x) = -0.005x2 + 0.25x, where h is the height in meters and x is the horizontal distance that the ball travels in meters. What is the maximum height that the ball reaches? ___________ m. Round your answer to the nearest thousandth.

OpenStudy (anonymous):

h(-.25/2(-.005))=h(25)

OpenStudy (anonymous):

you can also do it with calculus

OpenStudy (anonymous):

h(x) = -0.005x2 + 0.25x, h'(x)= -0.005(2x) +0.25 h'(x) =0 -0.005(2x) +0.25 =0 x=25 max height =h(25)=-0.005(25)^2 + 0.25(25)=3.125

OpenStudy (anonymous):

Jamen, are you in calculus, or algebra?

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

ok, do you know how to find the vertex of a parabola?

OpenStudy (anonymous):

not really im having trouble setting up and the equation and solving it..

OpenStudy (anonymous):

Given h(x) = -0.005x2 + 0.25x diff it once and set it to zero, we can get x that belongs to the vertex (slope of vertex=0) as the coeff. of x^2 <0, h(x) opens downward so we can know that the vertex is the max point put the x we get into h(x) to find its value

OpenStudy (anonymous):

The highway mileage m in miles per gallon for a compact car is approximated by m(s) = -0.025s2 + 2.45s - 30, where s is the speed in miles per hour. What is the maximum mileage for this compact car to the nearest tenth of a mile per gallon? __________ mi/gal. (Enter the number only.) can u also help me solve this equasion??

OpenStudy (anonymous):

The highway mileage m in miles per gallon for a compact car is approximated by m(s) = -0.025s2 + 2.45s - 30, where s is the speed in miles per hour. What is the maximum mileage for this compact car to the nearest tenth of a mile per gallon? __________ mi/gal. (Enter the number only.) can u also help me solve this equasion??

OpenStudy (anonymous):

CoCo, he said he's in algebra, not calculus

OpenStudy (anonymous):

Sorry for that. I study both so I cannot really distinguish which topic is belong to which side. Sorry again

OpenStudy (anonymous):

jamens you are familiar with quadratics, right? and can identify the values of a,b,c in the expression ax^2+bx+c?

OpenStudy (anonymous):

-0.025s^2+2.45s-30 Identify a,b,c

OpenStudy (anonymous):

ok i got -2.45/2(-0.025) -30 how would i solve it

OpenStudy (anonymous):

the x-coordinate of the vertex is -2.45/[2(-0.025)]

OpenStudy (anonymous):

when you plug the result of this into h(x), you will get your max. note that it is only a max because the coefficient of a<0 if a>0, plugging in this value would give you a min

OpenStudy (anonymous):

so the maximum mileage is m(49)

OpenStudy (anonymous):

plug 49 into m(s)

OpenStudy (anonymous):

i got 30.025 is that correct

OpenStudy (anonymous):

-60.025+120.05-30=30.025??

OpenStudy (anonymous):

30.025 is correct

OpenStudy (anonymous):

i typed it in and i keeps saying its incorrect..i have online school, so if i round that to the nearest tenth, would it be 30

OpenStudy (anonymous):

30.025 to the nearest tenth is 30.0

OpenStudy (anonymous):

oh i understand i jus typed in 30 instead of 30.0..thanks for your help